OpenStudy (anonymous):

While a planet P rotates in a circle about its sun, a moon M rotates in a circle about the planet, and both motions are in a plane. Let's call the distance between M and P one Suppose the distance of P from the sun is 3.5×103 lunar units; the planet makes one revolution about the sun every 9 years, and the moon makes one rotation about the planet every 0.25 years. Choosing coordinates centered at the sun, so that, at time t=0 the planet is at (3.5×103, 0), and the moon is at (3.5×103, 1), then the location of the moon at time t, where t is measured in years, is (x(t),y(t)), where x(t)= y(t)=

OpenStudy (anonymous):

Assuming 3.5x103 is: \[3.5 x 10^3\] To start, lets calculate the rotation of the planet around the sun. The period is \[(2\pi/9)t\] So for the x and y coordinate of the planet we get: \[x(t) = 3.5*10^3 \cos((2\pi/9)t)\] ; \[y(t) = 3.5*10^3 \sin((2\pi/9)t)\] For the moon, lets pretend the planet is stationary. The period for the moon around the planet is: \[(2\pi/.25)t = 8\pi t\] The x and y equations for the moon around the planet are: \[x(t) = \cos(8\pi t) ; y(t) = \sin(8\pi t)\] So, the complete equation for the motion of the moon around the sun, is the sum of these. We get: \[x(t) = 3.5*10^3 \cos((2\pi/9)t) + cos(8\pi t)\] \[y(t) = 3.5*10^3 \sin((2\pi/9)t) + sin(8\pi t)\]

OpenStudy (anonymous):

says they r both incorrect !